# MMHOI: Modeling Complex 3D Multi-Human Multi-Object Interactions

### v1

@@ Abstract
Building on MMHOI, we present MMHOI-Net, an end-to-end transformer-based neural network for jointly estimating human–object 3D geometries, their interactions, and associated actions. A key innovation in our framework is a structured dual-patch representation for modeling objects and their interactions, combined with action recognition to enhance the interaction prediction. Experiments on MMHOI and the recently proposed CORE4D datasets demonstrate that our approach achieves state-of-the-art performance in multi-HOI modeling, excelling in both accuracy and reconstruction quality.
@@ 1 Introduction
Perceiving 3D human-object interactions (HOIs) is central to understanding human behavior and is crucial for applications like action recognition, 3D scene reconstruction, and human-centric AI. While progress has been made in modeling 3D interactions between a single human and a single object [2, 5, 7, 30, 8, 17, 35, 3], real-world scenarios often involve multiple humans interacting with multiple objects simultaneously. These settings introduce greater complexity due to increased interaction patterns, diverse object categories, and ambiguity in recognizing multi-entity actions. Addressing these challenges is essential for developing robust models that generalize to the dynamic, collaborative nature of real-world environments.
In response, recent efforts have begun exploring more complex interaction settings beyond single-human scenarios. One such effort is HOI-M3 [36], which expands the scope to multiple humans and objects. However, it treats humans and objects as independent entities, overlooking the collaborative dynamics of real-world interactions. Recently, CORE4D [13] complements this by capturing 3D interactions with cooperative intent, but focuses only on two-person-single-object scenarios with limited action diversity. Moreover, the lack of varied action and detailed body part annotations limits these datasets’ ability to provide high-level task-specific guidance for low-level interaction modeling, leaving the synergy between 3D reconstruction and action recognition in multi-HOI scenarios largely unexplored.
@@
Building upon the MMHOI dataset, we propose a novel single-shot framework that simultaneously estimates the pose, shape, and 3D location of all humans and objects in camera space from a single RGB image. Human meshes are parameterized using the SMPL-X model [22], while object poses are represented with 6DoF parameters. Our framework leverages a Vision Transformer (ViT)-based architecture [4, 1] to extract patch-level features for both human and object modeling. To capture object-specific spatial and interaction cues, we introduce a structured dual-patch representation: each object is described by a main patch covering its primary region and a sub-patch encoding interaction-relevant features. This representation enables the model to reason about both object properties and fine-grained human-object interactions more effectively.
@@
  • The effectiveness of MMHOI-Net is demonstrated through extensive experiments on MMHOI, and also on the recent CORE4D [13] datasets.
@@ 2 Related Works
3D Single Human and Object Interaction. Several works that focusing on full-body interactions have been proposed to advance the HOI modeling [2, 8, 17, 30, 35, 3]. For instance, the BEHAVE dataset [2] benchmarks full-body human interactions with a single movable object, facilitating progress in modeling isolated HOI tasks [33, 20]. However, these datasets predominantly focus on interactions involving a single human and a single object, overlooking the more complex scenarios where multi-human interact with multi-object in cooperative environments.
3D Multiple Humans and Objects Interaction. Recently, the HOI-M3 dataset [36] advanced the field by capturing interactions among multi-humans and multi-objects in shared contextual environments. However, it lacks scenarios in which multiple humans simultaneously interact with one or more objects, leaving a significant gap for the explicit studying of cooperative interactions. More recently, CORE4D [13] focuses on two-person interactions involving a single object, capturing aspects of cooperation however has limited action annotations and do not consider multi-object scenarios. The absence of detailed action and interactive body part annotations in both datasets limits their scope in modeling interaction consistency and physical plausibility. Instead, our MMHOI dataset captures rich, complex interactions involving co-existing multi-human and multi-object-object interactions, with diverse action labels.
Action Recognition and 3D HOI. Action recognition and 3D human pose estimation are deeply interconnected. Early works primarily relied on 2D keypoints for action recognition [11, 32], however recent research has demonstrated that 3D pose estimation provides rich geometric information and context for complex interaction reasoning [15]. Although sparse 3D body joint representations often fail to capture fine-grained details, detailed 3D poses [23, 12] are seen to improve action recognition and single HOI modeling. Leveraging these insights, we propose to use action recognition to help enforce consistency in multi-HOI 3D estimation, ensuring that the reconstructed body poses remain faithful to their interaction objects.
@@ 3.1 Dataset Statistics
Our dataset captures realistic and diverse human-object interactions in 12 daily scenes, such as dining, collaborative work, and recreational activities (see Fig. 2). It comprises $\sim$600k frames featuring interactions performed by 13 participants (7 males and 6 females) with 22 commonly used objects spanning various categories. To complement the 3D and object annotations, MMHOI includes action annotations covering 78 distinct action classes and $14$ interacting body parts. Notably, MMHOI is the first dataset to combine 3D annotations of multi-human and multi-object-object with detailed action annotations, capturing their interactions in diverse and realistic scenarios. [^1] Each person appearing in our data set had to provide their consent, followed by obtaining the required IRB approvals for the use of human data in our study.
@@ 3.2 Dataset Acquisition
The MMHOI data was captured using a multi-camera setup consisting of four Microsoft Azure Kinect sensors. Each sensor recorded scenes at a resolution of 2048 $\times$ 1536 pixels and 30 fps. Calibration across the four sensors was performed using AprilTag markers [29] to ensure precise spatial alignment. Although the Azure Kinect cameras provide built-in synchronization, occasional data drift due to hardware limitations necessitated manual adjustments to achieve accurate temporal alignment.
@@ 3.3 Data Processing
Using synchronized and calibrated four-view videos, human segmentation was performed with the L-SAM [14]. We employ SMPL-X [22] as the 3D parametric human model to represent the statistical variations in human shape and pose. Following [1], the model parameters include pose parameters $\theta\in\mathbb{R}^{53\times 3}$ and shape parameters $\beta\in\mathbb{R}^{10}$. From these parameters, SMPL instantiates a human body mesh $M\in R^{N\times 3}$ with $N=10,475$ vertices. To initialize 3D human body representations from each camera view, we employed Multi-HMR [1], a single-view model that estimates SMPL-X parameters independently for each viewpoint. Professional annotators reviewed the results, selecting the view with minimal occlusion for refinement. These models were further refined using Iterative Closest Point (ICP) alignment with depth data (recorded from Kinect) in 3D space, followed by manual adjustments with FreeCAD [25] to finalize their positions.
@@
Object segmentation was also performed using L-SAM [14]. All objects in MMHOI are pre-scanned with the FreeScan UE Pro2 to obtain high-fidelity CAD models. These are initially aligned to depth maps (recorded from Kinect) via ICP to estimate 6D poses, then manually refined by annotators for accurate placement. We set the center of four Kinects as the world origin, and the CAD model’s center as the object’s origin. The object’s final pose ($R$, $T$) is computed via ICP between the original CAD template and the manually annotated placement.
@@
In addition to 3D annotations of humans and objects, MMHOI provides detailed action annotations to contextualize interactions. Each frame is labeled by trained annotators, who assign predefined action classes and specify which of the 14 body parts are involved in the interaction. The annotations are inspired by daily scenarios in video-based action datasets [6, 9, 18, 19, 21, 24, 26, 28, 31, 34, 38], and reflect the concept of synergy—how human actions and object affordances complement each other. Actions include both cooperative and individual tasks, such as “passing an object,” “assembling,” and “organizing.” Annotators cross-referenced multi-view RGB-D recordings to ensure accurate alignment between actions and 3D spatial configurations. Action labels are formatted as time-aligned sequences, enabling seamless integration with 3D and semantic data.
@@
GRAB [30]
BEHAVE [2]
CHAIRS [8]
NeuralDome [35]
InterCap [7]
HIMO [16]
ParaHome [10] ✓*
HOI-$\mathrm{M}^{3}$ [36] ✓* $\geq$3
CORE4D [13] 1 5 2
@@ 4 MMHOI-Net
This section introduces our MMHOI-Net, a vision-transformer based single-shot multi-task model for jointly learning the 3D human-object configurations, inferring the actions, and refining the body-part interactions. In Fig. 3, we illustrates the overall architecture of MMHOI-Net. The input to our model is a single RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ of resolution $H\times W$ depicting a multi-HOIs scene and a mask of this scene produced using L-SAM [14]. Our model outputs the SMPL-X parameters of all individual humans in $\mathbf{I}$, all the object meshes in the scene, their corresponding root 3D locations $t\in\mathbb{R}^{3}$ in camera coordinates, and associated action classes between all humans and object pairs.
@@ 4.1 Vision Transformer-Based Feature Encoding
To effectively process the input image, we employ a Vision Transformer (ViT) backbone [4] to extract high-level image features. The image $\mathbf{I}$ is first subdivided into non-overlapping patches of size $P\times P$. Each patch is embedded into a feature token, and these tokens are processed with self-attention blocks into $\mathbf{E}\in\mathbb{R}^{H/P\times W/P\times D}$ with dimension $D$ features. Similar to Multi-HMR [1], each output token maps spatially to a patch of the input RGB image. We employ the Human Perception Head (HPH) [1] for modeling the humans in the scene, and propose an Object Perception Head (OPH) for modeling the mesh parameters of the objects. Both these heads use the same ViT backbone, which is initialized with weights from a pre-trained MultiHMR model [1] during training. Below, we provide more details of HPH and our novel OPH. As depicted in Fig. 3, we further concatenate features from HPH and OPH for the action and body part classification.
Figure 3: MMHOI-Net model architecture. Given a single RGB image, our model jointly estimates the 3D geometry of multiple humans and objects while incorporating action recognition as a supervisory signal. A ViT backbone extracts patch-level features. For human perception, detected keypoints serve as queries in a Human Perception Head [1], regressing SMPL-X pose, shape, and translation. Object perception head uses a structured dual-patch representation to regress object 6DoF, center, and depth. An action MLP predicts action and interaction body part classes.
@@ 4.2 Human Perception Head
Following Multi-HMR [1], this head detects all primary human keypoints and predicts their SMPL-X parameters and depth. For $N$ humans in an image, a set of queries $\{\mathbf{q}_{n}\}_{n\in[N]}$ is initialized from the ViT feature tensor and processed in parallel through a stack of $L$ blocks ($L=2$ in the experiments), where alternating cross-attention and self-attention layers to refine the queries. The final output, $\mathbf{Q}^{L}\in\mathbb{R}^{(D+D^{\prime})\times N}$, is a set of $N$ refined human features used to regress the final human parameters (see [1] for details).
@@ 4.3 Object Perception Head
To jointly regress human and object geometries, we introduce an Object Perception Head (OPH) in our model (Fig. 3). Similar to the HPH branch for humans, OPH for objects follow a similar structure, however incorporates two patches instead of one – a main patch and a sub-patch – to model objects at multiple scales, which we call a structured dual-patch representation. Why do we need two patches? Recall that in [1], the human head is used as the primary keypoint to place the reconstructed human mesh. Objects in our dataset, however, vary in shape and symmetry, and scale. This is further exacerbated across categories, and thus treating them as monolithic entities is inadequate. To overcome this, we propose to use two patches overlapping the object mask to more robustly capture the object’s location, pose, and interaction points.
Our dual patch representation is obtained as follows. Given an object mask detected using L-SAM [14], the main patch for the object is defined as a patch covering the largest region of the mask containing the object mask center. The sub-patch is selected from one of the 8 surrounding patches to the main patch, corresponding to the second-largest region of the object mask and containing human-object interaction masks, i.e., patches containing mask parts belonging to both humans and the object (see Fig. 4). [^2] If no such interaction patches are found, we use the second largest patch containing the object mask.
@@
Similar to the processing pipeline in the HPH head, given $M$ objects in an image (detected by L-SAM), the final output of the transformer cross-attention layers is given by $\mathbf{Q}^{2\times L}\in\mathbb{R}^{2\times(D+D^{\prime}\times M)}$ and viewed as a set of $2\times M$ output object features. Next, the updated main and sub-patch queries from OPH are concatenated, and regressed using an MLP to produce the 6D object pose, center, and depth in 3D. To further enhance localization accuracy, available camera parameters are embedded into each patch as in [1].
@@ 4.4 Human-Object Interaction (Action) Head
We observe that action cues offer valuable semantic context, revealing the underlying structure of 3D human-object interactions. For instance, if an action module predicts “picking up a box”, the corresponding body parts (e.g., hands) must be touching the box. Inspired by prior works [20, 33] that use interactive body parts for 3D reconstruction in single-HOI, we extend this to multi-HOI by introducing an HOI head for action detection (see Fig. 3). To this end, we do a pairwise concatenation of the set of $N$ human features output from HPH and a set of $2\times M$ object features output from OPH, thus producing features $G=\{H_{main}\}_{N}\otimes\{(O_{main}\oplus O_{sub})\}_{M}$, where $\{H_{main}\}_{N}$ denotes the set of $N$ main patch (head) output features from the HPH branch and $\{(O_{main}\oplus O_{sub})\}_{M}$ indicate the concatenation of the output object main $O_{main}$ and sub-patch $O_{sub}$ features from OPH, respectively. After this step, we obtain $|G|=N\times M$ human-object pairs from the direct product, which is passed to a 2-layer MLP to jointly predict both action and interactive body part classes.

### v3

@@ Abstract
Building on MMHOI, we present MMHOI-Net, an end-to-end transformer-based neural network for jointly estimating human–object 3D geometries, their interactions, and associated actions. A key innovation in our framework is a structured dual-patch representation for modeling objects and their interactions, combined with action recognition to enhance the interaction prediction. Experiments on MMHOI and the recently proposed CORE4D datasets demonstrate that our approach achieves state-of-the-art performance in multi-HOI modeling, excelling in both accuracy and reconstruction quality. The MMHOI dataset is available at https://zenodo.org/records/17711786.
@@ 1 Introduction
Perceiving 3D human-object interactions (HOIs) is central to understanding human behavior and is crucial for applications like action recognition, 3D scene reconstruction, and human-centric AI. While progress has been made in modeling 3D interactions between a single human and a single object [bhatnagar2022behave, fan2023arctic, huang2024intercap, GRAB:2020, jiang2023chairs, mandery2016unifying, zhang2023neuraldome, cseke_tripathi_2025_pico], real-world scenarios often involve multiple humans interacting with multiple objects simultaneously. These settings introduce greater complexity due to increased interaction patterns, diverse object categories, and ambiguity in recognizing multi-entity actions. Addressing these challenges is essential for developing robust models that generalize to the dynamic, collaborative nature of real-world environments.
In response, recent efforts have begun exploring more complex interaction settings beyond single-human scenarios. One such effort is HOI-M3 [zhang2024hoi], which expands the scope to multiple humans and objects. However, it treats humans and objects as independent entities, overlooking the collaborative dynamics of real-world interactions. Recently, CORE4D [liu2025core4d] complements this by capturing 3D interactions with cooperative intent, but focuses only on two-person-single-object scenarios with limited action diversity. Moreover, the lack of varied action and detailed body part annotations limits these datasets’ ability to provide high-level task-specific guidance for low-level interaction modeling, leaving the synergy between 3D reconstruction and action recognition in multi-HOI scenarios largely unexplored.
@@
Building upon the MMHOI dataset, we propose a novel single-shot framework that simultaneously estimates the pose, shape, and 3D location of all humans and objects in camera space from a single RGB image. Human meshes are parameterized using the SMPL-X model [SMPL-X:2019], while object poses are represented with 6DoF parameters. Our framework leverages a Vision Transformer (ViT)-based architecture [dosovitskiy2020vit, multi-hmr2024] to extract patch-level features for both human and object modeling. To capture object-specific spatial and interaction cues, we introduce a structured dual-patch representation: each object is described by a main patch covering its primary region and a sub-patch encoding interaction-relevant features. This representation enables the model to reason about both object properties and fine-grained human-object interactions more effectively.
@@
  • The effectiveness of MMHOI-Net is demonstrated through extensive experiments on MMHOI, and also on the recent CORE4D [liu2025core4d] datasets.
@@ 2 Related Works
3D Single Human and Object Interaction. Several works that focusing on full-body interactions have been proposed to advance the HOI modeling [bhatnagar2022behave, jiang2023chairs, mandery2016unifying, GRAB:2020, zhang2023neuraldome, cseke_tripathi_2025_pico]. For instance, the BEHAVE dataset [bhatnagar2022behave] benchmarks full-body human interactions with a single movable object, facilitating progress in modeling isolated HOI tasks [xie2022chore, nam2024joint]. However, these datasets predominantly focus on interactions involving a single human and a single object, overlooking the more complex scenarios where multi-human interact with multi-object in cooperative environments.
3D Multiple Humans and Objects Interaction. Recently, the HOI-M3 dataset [zhang2024hoi] advanced the field by capturing interactions among multi-humans and multi-objects in shared contextual environments. However, it lacks scenarios in which multiple humans simultaneously interact with one or more objects, leaving a significant gap for the explicit studying of cooperative interactions. More recently, CORE4D [liu2025core4d] focuses on two-person interactions involving a single object, capturing aspects of cooperation however has limited action annotations and do not consider multi-object scenarios. The absence of detailed action and interactive body part annotations in both datasets limits their scope in modeling interaction consistency and physical plausibility. Instead, our MMHOI dataset captures rich, complex interactions involving co-existing multi-human and multi-object-object interactions, with diverse action labels.
Action Recognition and 3D HOI. Action recognition and 3D human pose estimation are deeply interconnected. Early works primarily relied on 2D keypoints for action recognition [TIN-net, PMF], however recent research has demonstrated that 3D pose estimation provides rich geometric information and context for complex interaction reasoning [luvizon2020multi]. Although sparse 3D body joint representations often fail to capture fine-grained details, detailed 3D poses [rajasegaran2023benefits, li2020detailed] are seen to improve action recognition and single HOI modeling. Leveraging these insights, we propose to use action recognition to help enforce consistency in multi-HOI 3D estimation, ensuring that the reconstructed body poses remain faithful to their interaction objects.
@@ 3.1 Dataset Statistics
Our dataset captures realistic and diverse human-object interactions in 12 daily scenes, such as dining, collaborative work, and recreational activities (see Fig. 2). It comprises $\sim$600k frames featuring interactions performed by 13 participants (7 males and 6 females) with 22 commonly used objects spanning various categories. To complement the 3D and object annotations, MMHOI includes action annotations covering 78 distinct action classes and $14$ interacting body parts. Notably, MMHOI is the first dataset to combine 3D annotations of multi-human and multi-object-object with detailed action annotations, capturing their interactions in diverse and realistic scenarios. [^1] Each subject in our dataset provided their written consent to be included.
@@ 3.2 Dataset Acquisition
The MMHOI data was captured using a multi-camera setup consisting of four Microsoft Azure Kinect sensors. Each sensor recorded scenes at a resolution of 2048 $\times$ 1536 pixels and 30 fps. Calibration across the four sensors was performed using AprilTag markers [song7extrinsic] to ensure precise spatial alignment. Although the Azure Kinect cameras provide built-in synchronization, occasional data drift due to hardware limitations necessitated manual adjustments to achieve accurate temporal alignment.
@@ 3.3 Data Processing
Using synchronized and calibrated four-view videos, human segmentation was performed with the L-SAM [lsam]. We employ SMPL-X [SMPL-X:2019] as the 3D parametric human model to represent the statistical variations in human shape and pose. Following [multi-hmr2024], the model parameters include pose parameters $\theta\in\mathbb{R}^{53\times 3}$ and shape parameters $\beta\in\mathbb{R}^{10}$. From these parameters, SMPL instantiates a human body mesh $M\in R^{N\times 3}$ with $N=10,475$ vertices. To initialize 3D human body representations from each camera view, we employed Multi-HMR [multi-hmr2024], a single-view model that estimates SMPL-X parameters independently for each viewpoint. Professional annotators reviewed the results, selecting the view with minimal occlusion for refinement. These models were further refined using Iterative Closest Point (ICP) alignment with depth data (recorded from Kinect) in 3D space, followed by manual adjustments with FreeCAD [riegel2016freecad] to finalize their positions.
@@
Object segmentation was also performed using L-SAM [lsam]. All objects in MMHOI are pre-scanned with the FreeScan UE Pro2 to obtain high-fidelity CAD models. These are initially aligned to depth maps (recorded from Kinect) via ICP to estimate 6D poses, then manually refined by annotators for accurate placement. We set the center of four Kinects as the world origin, and the CAD model’s center as the object’s origin. The object’s final pose ($R$, $T$) is computed via ICP between the original CAD template and the manually annotated placement.
@@
In addition to 3D annotations of humans and objects, MMHOI provides detailed action annotations to contextualize interactions. Each frame is labeled by trained annotators, who assign predefined action classes and specify which of the 14 body parts are involved in the interaction. The annotations are inspired by daily scenarios in video-based action datasets [AVA, CMU_Panoptic, Hollywood2, Moments, TV_Human_Interaction, SALSA, UT-Interaction, Kinetics700, shakefive, SBU_Kinect, HACS], and reflect the concept of synergy—how human actions and object affordances complement each other. Actions include both cooperative and individual tasks, such as “passing an object,” “assembling,” and “organizing.” Annotators cross-referenced multi-view RGB-D recordings to ensure accurate alignment between actions and 3D spatial configurations. Action labels are formatted as time-aligned sequences, enabling seamless integration with 3D and semantic data.
@@
GRAB [GRAB:2020]
BEHAVE [bhatnagar2022behave]
CHAIRS [jiang2023chairs]
NeuralDome [zhang2023neuraldome]
InterCap [huang2024intercap]
HIMO [lv2024himonewbenchmarkfullbody]
ParaHome [kim2024parahome] ✓*
HOI-$\mathrm{M}^{3}$ [zhang2024hoi] ✓* $\geq$3
CORE4D [liu2025core4d] 1 5 2
@@ 4 MMHOI-Net
This section introduces our MMHOI-Net, a vision-transformer based single-shot multi-task model for jointly learning the 3D human-object configurations, inferring the actions, and refining the body-part interactions. In Fig. 3, we illustrates the overall architecture of MMHOI-Net. The input to our model is a single RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ of resolution $H\times W$ depicting a multi-HOIs scene and a mask of this scene produced using L-SAM [lsam]. Our model outputs the SMPL-X parameters of all individual humans in $\mathbf{I}$, all the object meshes in the scene, their corresponding root 3D locations $t\in\mathbb{R}^{3}$ in camera coordinates, and associated action classes between all humans and object pairs.
@@ 4.1 Vision Transformer-Based Feature Encoding
To effectively process the input image, we employ a Vision Transformer (ViT) backbone [dosovitskiy2020vit] to extract high-level image features. The image $\mathbf{I}$ is first subdivided into non-overlapping patches of size $P\times P$. Each patch is embedded into a feature token, and these tokens are processed with self-attention blocks into $\mathbf{E}\in\mathbb{R}^{H/P\times W/P\times D}$ with dimension $D$ features. Similar to Multi-HMR [multi-hmr2024], each output token maps spatially to a patch of the input RGB image. We employ the Human Perception Head (HPH) [multi-hmr2024] for modeling the humans in the scene, and propose an Object Perception Head (OPH) for modeling the mesh parameters of the objects. Both these heads use the same ViT backbone, which is initialized with weights from a pre-trained MultiHMR model [multi-hmr2024] during training. Below, we provide more details of HPH and our novel OPH. As depicted in Fig. 3, we further concatenate features from HPH and OPH for the action and body part classification.
Figure 3: MMHOI-Net model architecture. Given a single RGB image, our model jointly estimates the 3D geometry of multiple humans and objects while incorporating action recognition as a supervisory signal. A ViT backbone extracts patch-level features. For human perception, detected keypoints serve as queries in a Human Perception Head [multi-hmr2024], regressing SMPL-X pose, shape, and translation. Object perception head uses a structured dual-patch representation to regress object 6DoF, center, and depth. An action MLP predicts action and interaction body part classes.
@@ 4.2 Human Perception Head
Following Multi-HMR [multi-hmr2024], this head identifies all major human keypoints and predicts their SMPL-X parameters and depth. For $N$ humans in an image, a set of queries $\{\mathbf{q}_{n}\}_{n\in[N]}$ is initialized from the ViT feature tensor and processed in parallel through a stack of $L$ blocks ($L=2$ in the experiments), where alternating cross-attention and self-attention layers to progressively refine the queries. The final output, $\mathbf{Q}^{L}\in\mathbb{R}^{(D+D^{\prime})\times N}$, is a set of $N$ refined human features used to regress the final human parameters (see [multi-hmr2024] for further details).
@@ 4.3 Object Perception Head
To jointly regress human and object geometries, we introduce an Object Perception Head (OPH) in our model (Fig. 3). Similar to the HPH branch for humans, OPH for objects follow a similar structure, however incorporates two patches instead of one – a main patch and a sub-patch – to model objects at multiple scales, which we call a structured dual-patch representation. Why do we need two patches? Recall that in [multi-hmr2024], the human head is used as the primary keypoint to place the reconstructed human mesh. Objects in our dataset, however, vary in shape and symmetry, and scale. This is further exacerbated across categories, and thus treating them as monolithic entities is inadequate. To overcome this, we propose to use two patches overlapping the object mask to more robustly capture the object’s location, pose, and interaction points.
Our dual patch representation is obtained as follows. Given an object mask detected using L-SAM [lsam], the main patch for the object is defined as a patch covering the largest region of the mask containing the object mask center. The sub-patch is selected from one of the 8 surrounding patches to the main patch, corresponding to the second-largest region of the object mask and containing human-object interaction masks, i.e., patches containing mask parts belonging to both humans and the object (see Fig. 4). [^2] If no such interaction patches are found, we use the second largest patch containing the object mask.
@@
Similar to the processing pipeline in the HPH head, given $M$ objects in an image (detected by L-SAM), the final output of the transformer cross-attention layers is given by $\mathbf{Q}^{2\times L}\in\mathbb{R}^{2\times(D+D^{\prime}\times M)}$ and viewed as a set of $2\times M$ output object features. Next, the updated main and sub-patch queries from OPH are concatenated, and regressed using an MLP to produce the 6D object pose, center, and depth in 3D. To further enhance localization accuracy, available camera parameters are embedded into each patch as in [multi-hmr2024].
@@ 4.4 Human-Object Interaction (Action) Head
We observe that action cues offer valuable semantic context, revealing the underlying structure of 3D human-object interactions. For instance, if an action module predicts “picking up a box”, the corresponding body parts (e.g., hands) must be touching the box. Inspired by prior works [nam2024joint, xie2022chore] that use interactive body parts for 3D reconstruction in single-HOI, we extend this to multi-HOI by introducing an HOI head for action detection (see Fig. 3). To this end, we do a pairwise concatenation of the set of $N$ human features output from HPH and a set of $2\times M$ object features output from OPH, thus producing features $G=\{H_{main}\}_{N}\otimes\{(O_{main}\oplus O_{sub})\}_{M}$, where $\{H_{main}\}_{N}$ denotes the set of $N$ main patch (head) output features from the HPH branch and $\{(O_{main}\oplus O_{sub})\}_{M}$ indicate the concatenation of the output object main $O_{main}$ and sub-patch $O_{sub}$ features from OPH, respectively. After this step, we obtain $|G|=N\times M$ human-object pairs from the direct product, which is passed to a 2-layer MLP to jointly predict both action and interactive body part classes.